# MetaAligner: Towards Generalizable Multi-Objective Alignment of Language Models

### v2

@@ 3.3 Model Training
Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs. However, equal preferences are common in many scenarios [34, 8] and enclose useful information such as the principle components of preference modeling regarding each objective. Based on these intuitions, we introduce a novel equal-preference modeling stage to fine-tune the warmed-up MetaAligner on the equal-preference subset $\mathcal{D}_{e}$.
Step 3: preference alignment
This stage fine-tunes the MetaAligner on the preference subset $\mathcal{D}_{p}$, which instructs the model to perform conditional weak-to-strong correction on the dis-preferred response and achieve alignment to the preferred response on the specified objectives.
@@ 3.4 Generalizable Inference
[…]
This simple pattern can theoretically lead to unlimited simultaneous alignment objectives. We expect MetaAligner to make generalizable weak-to-strong corrections under these unseen conditions via its in-context learning ability. This advancement marks a new exploration into generalizable multi-objective preference alignment.
@@ 4.1 Experimental Settings
We train and evaluate on 3 alignment datasets: (1) HH-RLHF [4]: a large-scale dataset with 160K prompts and corresponding response pairs. We follow previous works [35] and utilize 3 open-sourced reward models on "Harmless", "Helpful", and "Humor" to assess and rank the responses in their corresponding objectives; (2) UltraFeedback [8]: a multi-aspect alignment dataset with 64K prompts with preferences obtained from GPT-4. We select "Instruction following", "Honest", "Truthful", and "Helpful" objectives during training; (3) IMHI: we evaluate MetaAligner on domain knowledge-intense scenarios by creating another alignment dataset based on the IMHI dataset [34] targeting interpretable mental health analysis, which consists of 9 mental health analysis sub-tasks. We invite domain experts to manually label 7.2K response pairs sampled from the training split of IMHI, considering 3 objectives: "Correct", "Informative", and "Professional".
@@
We train MetaAligner-(1.1B, 7B, 13B) models based on TinyLLaMA-1.1B [36] and LLaMA2-(7B, 13B) [31] foundation models. We utilize MetaAligner to perform multi-objective alignment on the following open-source policy models: LLaMA2-Chat-(7B,13B,70B) [31], Gemma-instruct-(2B,7B) [30], and Vicuna-(7B, 13B, 33B) [7]. We also align 2 advanced close-sourced foundation models: GPT-3.5 [22] and Claude-3-Sonnet [3], where model parameters are inaccessible.
@@
On each objective, we quantify the alignment performance of model outputs by comparing their win rates against the ground-truth response provided by the benchmark datasets. Considering the large amounts of test samples, we utilize GPT-4, a widely utilized evaluation tool in previous works [11, 29, 19], to perform the judgments. Each target response, ground-truth response, query, and evaluated objectives are provided via prompt engineering. GPT-4 is required to compare and select the response with higher alignment to the specified objectives or indicate a tied performance.
More details about the training process, model cards, dataset statistics, IMHI dataset annotation, and evaluation settings are presented in Appendix C.
@@
13B LLaMA2-Chat-7B +21.0% +30.5% +23.75% +51.83% +47.5% +45.33% +38.67% +28.33% +38.33% +50.56%
LLaMA2-Chat-13B +22.75% +16.08% +12.25% +46.33% +48.67% +46.83% +41.17% +30.56% +37.22% +40.56%
LLaMA2-Chat-70B +16.58% +11.92% +23.08% +44.33% +35.0% +45.5% +24.0% +31.67% +30.56% +36.12%
Gemma-instruct-2B +27.5% +20.75% +24.33% +55.0% +44.67% +51.33% +36.83% +35.55% +63.33% +65.0%
Gemma-instruct-7B +14.0% +27.75% +25.17% +42.0% +40.17% +35.17% +31.17% +34.45% +50.0% +49.44%
Vicuna-7B +18.0% +12.83% +30.33% +41.5% +39.83% +44.33% +37.5% +24.44% +23.33% +21.11%
Vicuna-13B +17.92% +29.0% +32.67% +47.33% +49.17% +47.0% +40.67% +28.33% +23.34% +18.9%
Vicuna-33B +31.5% +17.09% +27.83% +50.5% +53.17% +45.83% +38.5% +23.89% +23.89% +14.45%
GPT-3.5 +19.92% +25.0% +23.84% +40.33% +40.17% +36.83% +23.67% +26.67% +25.66% +33.62%
Claude-3-Sonnet +27.17% +23.58% +27.17% +38.5% +39.5% +37.67% +29.83% +28.67% +20.0% +11.2%
Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on 3 datasets over different policy models. The responses are simultaneously aligned on all objectives, then tested on each objective. "IF" denotes the "Instruction following" objective. "+" shows the advantage of aligned outputs over the unaligned outputs on win rates against the golden responses.
@@ 4.2 Overall Performance
MetaAligner-(1.1B, 7B, 13B) performance on 3 alignment datasets are shown in Table 2. According to the results, the MetaAligner models achieve substantial improvement for most objectives and policy models. For example, on UltraFeedback, there is an average of 11.47% advantage for MetaAligner-1.1B on "Honest", 34.39% for MetaAligner-7B, and 43.79% for MetaAligner-13B. These results show the general effectiveness of MetaAligner on various upstream models and the feasibility of plug-and-play multi-objective alignment. On the mental health analysis benchmark IMHI, MetaAligner models also show remarkable win rates on all objectives, proving their effectiveness in performing multi-objective alignment in domain knowledge-intense scenarios. We further evaluate MetaAligner on each IMHI sub-task and the results are shown in Appendix D.
@@
From the policy model scale perspective, MetaAligner provides successful alignments to open-sourced models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself. In the extreme case, MetaAligner-1.1B advances the win rates of LLaMA2-Chat-70B outputs, a policy model with 63$\times$ more parameters, by an average of 12.19% on HH-RLHF, 13.08% on UltraFedback, and 13.26% on IMHI. These results prove MetaAligner as a parameter-efficient alignment strategy compared to previous multi-objective alignment methods, where the policy model weights are updated, leading to an inevitable surge of computation resources as policy model sizes grow. MetaAligner also significantly improves performance on API-based models: GPT-3.5 and Claude-3-Sonnet. These results prove its potential for application in close-sourced scenarios and effective multi-objective alignment of state-of-the-art policy models.
@@
Within most policy model families, we observe a decrease in win-rate advantage as their sizes increase. These decreases indicate a struggle aligning powerful large-scale models with a small MetaAligner. Fortunately, MetaAligner’s capabilities also show scalability. Increasing the size of its base model leads to a higher win-rate advantage on most policy models. For example, on UltraFeedback, MetaAligner-7B outperforms MetaAligner-1.1B on all 10 policy models, and MetaAligner-13B further surpasses MetaAligner-7B by an average of 12.58%. These observations motivate further explorations in model scale-performance balance for MetaAligner.
@@ 4.3 Evaluations of Objective-Wise Alignment
We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives. We utilize MetaAligner to perform 6 levels of alignments: unaligned, aligning on each objective (Instruction following, Helpfulness, Honesty, and Truthfulness), and full alignment. We leverage GPT-4 to score the responses ranging from 0 to 10. Experimental details and more results are shown in Appendix E. The results are shown in Figure 2. We have the following observations:
Objective-wise alignment improves performance on the target objective and benefits the performance on other objectives. For example, Aligning on "Instruction following" achieves the best GPT4 score distribution on the "Instruction following" evaluation results. It also significantly increases GPT4 scores on "Helpful", "Honest", and "Truthful" over the unaligned responses. This tendency holds with other policy models and alignment objectives. These results further prove the complex interplay within multi-objective alignment, where positive correlations and contradictions [11] co-exist.
@@
Figure 3: Zero-shot alignment on 6 unseen objectives based on LLaMA2-Chat-70B outputs of UltraFeedback. In the x-axis, "Aligned Obj." denotes the 4 supervised objectives ($\diamond$" markers), and "+" denotes further addition of an unseen objective ($\circ$" markers). $\star$" denotes the unaligned win rates for the unseen objectives, "-." lines identify win rate fluctuations before alignment, and solid lines identify fluctuations after alignment.
4.4 Generalizable Alignment
In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with 6 less relevant objectives that are unseen during its training phases: Specific, Factual, Readable, Fair, Repeat, and Length [10]. More details about these objectives are in Appendix B.
We randomly select 2,700 queries from the UltraFeedback dataset and re-align the LLaMA2-Chat-70B outputs with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one, with 10 aligned objectives in total. Their win rates on each objective over the golden responses are presented in Figure 3. According to the results, we have the following conclusions:
Performance on aligned objectives is maintained with more unseen alignment objectives. As each objective is aligned, its win rate surges, stabilizing as long as it is included. On simultaneously aligning 10 objectives, MetaAligner-7B outperforms LLaMA2-Chat-70B outputs by an average of 14.25% on unseen objectives. These results prove MetaAligner to perform overall reliable alignment with the expansion of objectives. However, aligning on new objectives can affect other objectives. For example, aligning on "Fair" with MetaAligner-(7B, 13B) benefits its win rates, but harms performance on objectives such as "Readable" and "Factual". The decreases are due to the "alignment tax" [11], where enhancements in one objective can diminish performance in other objectives due to their controversial nature.
MetaAligner’s generalizability shows scalability. Performance on the 6 unseen objectives increases with the scale-up of MetaAligner model size. MetaAligner-1.1B provides limited alignment on most unseen objectives, but MetaAligner-7B extends the win rates to an average of 48.5%, and MetaAligner-13B further reaches 61.25%. MetaAligner-13B also more effectively aligns objectives such as "Length", where smaller models perform badly. This scalability is attributed to larger foundation models’ growing in-context learning ability, which enables accurate interpretations of the objective descriptions and instructions. These observations motivate further explorations into the correlation between generalizable alignment and model scales in future work.
HH-RLHF UltraFeedback
Algorithm GPU Hours Harmless Helpful Humour Avg. IF Honest Truthful Helpful Avg.
MORLHF 1892.3 62.83% 51.2% 77.5% 63.84% 32.18% 33.7% 26.1% 33.7% 31.42%
MODPO 405.9 65.0% 64.0% 78.0% 69.0% 30.82% 43.4% 37.19% 25.0% 34.1%
SFT 247.34 66.5% 75.0% 76.5% 72.67% 27.0% 36.5% 26.0% 36.5% 31.5%
MetaAligner-13B 403.44 73.5% 85.5% 86.0% 81.67% 68.5% 59.5% 64.0% 55.0% 61.75%
Table 3: Comparisons of win rates between alignment methods based on the LLaMA2-Chat-7B policy model. "GPU hours" records the summed GPU running time on all datasets.
4.5 MetaAligner vs. MORLHF/MODPO/SFT-based Method
We compare the performance of MetaAligner with MORLHF, MODPO, and SFT-based methods on the HH-RLHF and UltraFeedback datasets, where explicit rewards are available. We implement the linear scalarization method for MORLHF, the CDPO [11] realization of MODPO, and RiC [35] realization of the SFT-based method. We select LLaMA2-Chat-7B as the policy model, and the results are presented in Table 3. Appendix F presents details about the baseline model implementations and GPU hours calculations.
According to the results, MetaAligner-13B significantly outperforms all other methods with an average of 81.67% win rate on HH-RLHF and 61.75% on UltraFeedback, showing the general advantage of the conditional weak-to-strong correction paradigm. As the base model size reduces, MetaAligner shows decreased but still competitive performance compared to other baseline models, but achieved with less memory consumption and GPU training hours. Impressively, the MetaAligner-1.1B model achieves comparable average performance to MORLHF, MODPO, and SFT-based methods on both datasets, but with 2.05$\times$-15.67$\times$ less GPU training hours and 6.36$\times$ smaller size than the LLaMA2-Chat-7B policy model. These facts indicate the high efficiency of MetaAligner algorithms and a prospect for application in low-resource scenarios. Compared to previous methods, MetaAligner models can also achieve balanced performances in objective-wise evaluations, including contradictory objectives such as "Harmless" and "Helpful", without requiring explicit hyper-parameter tuning for achieving Pareto optimal solutions [11, 35, 20]. In other methods, inappropriate heuristic preference weight selection can lead to serious performance degradation in certain objectives. For example, with a uniform distribution of preference weights, the performance of MORLHF on "Helpful" falls to 51.2%, a huge gap to other methods.

### v3

@@ 3.2.2 Three-Step Model Training
In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process. We propose a three-step paradigm based on the objective function in Eqn. 6: (1) Warming up. This stage trains the model in identical response pairs with a warm-up subset, a prelude proven effective in residual correction strategies [12, 13]. We randomly sample a subset of the equal subset $\mathcal{D}_{e}$ as the warm-up subset, but set an identical target response for each instance; (2) Equal-preference alignment. Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs. However, equal preferences are common in many scenarios [35, 8] and enclose useful information such as the principle components of preference modeling regarding each objective. Based on these intuitions, we introduce a novel equal-preference alignment step to fine-tune the warmed-up MetaAligner on the equal subset $\mathcal{D}_{e}$; (3) Contrastive-preference alignment. This stage fine-tunes the MetaAligner on the contrastive preference subset $\mathcal{D}_{c}$, which instructs the model to perform conditional weak-to-strong correction on the specified objectives.
@@ 4.1 Experimental Settings
We transfer the following three alignment datasets into dynamic multi-objective datasets: (1) HH-RLHF [4]: a large-scale dataset with 160K prompts and corresponding response pairs. We follow Yang et al. [36] and use open-sourced reward models on three objectives: "Harmless", "Helpful", and "Humor" to score the responses; (2) UltraFeedback [8]: a multi-aspect alignment dataset with 64K prompts with preferences obtained from GPT-4, including "Instruction following", "Honest", "Truthful", and "Helpful" objectives; (3) IMHI: we create an alignment dataset on the IMHI dataset [35] targeting interpretable mental health analysis. We invite domain experts to label 7.2K response pairs considering 3 objectives: "Correct", "Informative", and "Professional". Figure 2 shows the objective distributions on two datasets. The objectives display balanced overall distributions across objective set sizes, training MetaAligner to adjust targets dynamically. Most objectives also cover considerable proportions in each column category, alleviating label imbalance problems.
Figure 2: Heatmaps of the objective distributions. The columns categorize samples according to the sizes of their objective set. For the lines, "Overall" shows their distributions in the training data. Other lines show objective-wise distributions across different categories in the columns.
@@
We train MetaAligner-(1.1B, 7B, 13B) models based on TinyLLaMA-1.1B [38] and LLaMA2-(7B, 13B) [31] foundation models. We utilize MetaAligner to perform multi-objective alignment on the following open-source policy models: LLaMA2-Chat-(7B,13B,70B) [31], Gemma-instruct-(2B,7B) [30], and Vicuna-(7B, 13B, 33B) [7]. We also align two advanced close-source foundation models: GPT-3.5-Turbo [22] and Claude-3-Sonnet [3], where model parameters are inaccessible.
@@
On each objective, we quantify the alignment performance of model outputs by comparing their win rates against the ground-truth response provided by the benchmark datasets. Considering the large amounts of test samples, we leverage GPT-4 [2], a widely utilized evaluation tool in previous works [11, 29, 19], to perform the judgments. Each target response, ground-truth response, query, and evaluated objectives are provided via prompt engineering. GPT-4 is required to compare and select the response with higher alignment on the specified objective or indicate a tied performance of the two responses.
More details about the training process, model cards, dataset statistics, IMHI dataset annotation, and evaluation settings are presented in Appendix D.
Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on 3 datasets over different policy models. The responses are simultaneously aligned on all trained objectives, then evaluated on each objective. "IF" denotes the "Instruction following" objective. "+" shows the advantage of aligned outputs over the unaligned outputs on win rates against the ground-truth responses.
@@
13B LLaMA2-Chat-7B +24.0% +30.5% +23.75% +51.83% +47.5% +45.33% +38.67% +28.33% +38.33% +50.56%
LLaMA2-Chat-13B +17.75% +16.58% +15.75% +46.33% +48.67% +46.83% +41.17% +30.56% +37.22% +40.56%
LLaMA2-Chat-70B +16.58% +19.42% +26.58% +44.33% +35.0% +45.5% +24.0% +31.67% +30.56% +36.12%
Gemma-instruct-2B +18.5% +17.25% +24.33% +55.0% +44.67% +51.33% +36.83% +35.55% +63.33% +65.0%
Gemma-instruct-7B +17.5% +23.75% +30.17% +42.0% +40.17% +35.17% +31.17% +34.45% +50.0% +49.44%
Vicuna-7B +19.0% +19.83% +26.33% +41.5% +39.83% +44.33% +37.5% +24.44% +23.33% +21.11%
Vicuna-13B +18.92% +28.5% +32.67% +47.33% +49.17% +47.0% +40.67% +28.33% +23.34% +18.9%
Vicuna-33B +31.5% +20.09% +27.83% +50.5% +53.17% +45.83% +38.5% +23.89% +23.89% +14.45%
GPT-3.5-Turbo +18.42% +25.0% +29.34% +40.33% +40.17% +36.83% +23.67% +26.67% +25.66% +33.62%
Claude-3-Sonnet +21.17% +20.58% +27.17% +38.5% +39.5% +37.67% +29.83% +28.67% +20.0% +11.2%
@@ 4.2 Overall Performance
MetaAligner-(1.1B, 7B, 13B) performance on 3 alignment datasets are shown in Table 2. According to the results, the MetaAligner models achieve substantial improvement for most objectives and policy models. For example, on UltraFeedback, there is an average of 11.47% advantage for MetaAligner-1.1B on "Honest", 34.39% for MetaAligner-7B, and 43.79% for MetaAligner-13B. These results show the general effectiveness of MetaAligner on various upstream models and the feasibility of plug-and-play multi-objective alignment. On the mental health analysis benchmark IMHI, MetaAligner models also show remarkable win rates on all objectives, proving their effectiveness in performing multi-objective alignment in domain knowledge-intense scenarios. We further evaluate MetaAligner on each IMHI sub-task and the results are shown in Appendix E.
@@
From the policy model scale perspective, MetaAligner provides successful alignments to open-source models with sizes ranging from 2B to 70B, significantly extending the size of MetaAligner itself. In the extreme case, MetaAligner-1.1B advances the win rates of LLaMA2-Chat-70B outputs, a policy model with 63$\times$ more parameters, by an average of 12.19% on HH-RLHF, 13.08% on UltraFeedback, and 13.26% on IMHI. These results prove MetaAligner as a parameter-efficient alignment strategy compared to previous multi-objective alignment methods, where the policy model weights are updated, leading to an inevitable surge of computation resources as policy model sizes grow. MetaAligner also significantly improves performance on close-source LLMs: GPT-3.5-Turbo and Claude-3-Sonnet. These results prove its potential for application in close-source scenarios and effective multi-objective alignment of state-of-the-art policy models.
@@
Within most policy model families, we observe a decreasing trend in win-rate advantage as their sizes increase. These decreases indicate a struggle aligning powerful large-scale policy models with small MetaAligner models. Fortunately, MetaAligner’s capabilities also show scalability. Increasing the size of its base model leads to a higher win-rate advantage on most policy models. For example, on UltraFeedback, MetaAligner-7B outperforms MetaAligner-1.1B on all 10 policy models, and MetaAligner-13B further surpasses MetaAligner-7B by an average of 12.58%. These observations motivate further explorations in model size-performance balance for MetaAligner.
Table 3: Comparisons of win rates between alignment methods. "GPU Hours" records the summed GPU running time on all datasets. "-Equal Pref." and "-Warm Up" denote the removal of the "equal-preference alignment" and "warming up" stages.
HH-RLHF UltraFeedback
Policy Model Algorithm GPU Hours Harmless Helpful Humour Avg. IF Honest Truthful Helpful Avg.
LLaMA2-Chat-7B MORLHF 1892.3 62.83% 51.2% 77.5% 63.84% 32.18% 33.7% 26.1% 33.7% 31.42%
MODPO 405.9 65.0% 64.0% 78.0% 69.0% 30.82% 43.4% 37.19% 25.0% 34.1%
SFT 247.34 66.5% 75.0% 76.5% 72.67% 27.0% 36.5% 26.0% 36.5% 31.5%
Aligner-7B 236.8 72.0% 81.9% 70.12% 74.67% 52.38% 44.23% 37.19% 39.1% 43.23%
-Equal Pref. 73.82% 80.7% 77.39% 77.3% 46.8% 43.6% 53.17% 41.7% 46.32%
-Warm Up 77.1% 80.32% 82.63% 80.02% 49.96% 47.4% 55.73% 44.18% 49.32%
MetaAligner-13B 403.44 76.5% 85.5% 86.0% 82.67% 68.5% 59.5% 64.0% 55.0% 61.75%
LLaMA2-Chat-70B Self-Refinement 70.48% 82.8% 68.91% 74.06% 49.95% 62.91% 60.77% 57.6% 55.05%
MetaAligner-7B 242.68 85.16% 89.42% 88.08% 87.55% 67.05% 63.72% 70.1% 54.7% 63.89%
4.3 MetaAligner vs. Baseline Methods
We compare the performance of MetaAligner with MORLHF, MODPO, SFT-based methods, and Aligner. We implement the linear scalarization method for MORLHF, the CDPO [11] realization of MODPO, and RiC [36] realization of the SFT-based method. As Aligner is not suitable for multi-objective alignment, we train Aligner-7B on "Helpful" annotations for HH-RLHF and "IF" for UltraFeedback. We compare these methods on the LLaMA2-Chat-7B policy model. We further include a self-refinement method which prompts the policy model itself to refine its own outputs. We compare self-refinement on the LLaMA2-Chat-70B policy model as it requires strong in-context learning ability from the policy model. The results are presented in Table 3. Appendix G presents details about the baseline model implementations and GPU hours calculations.
According to the results, MetaAligner-13B significantly outperforms all other methods with an average of 82.67% win rate on HH-RLHF and 61.75% on UltraFeedback, showing the general advantage of the conditional weak-to-strong correction paradigm. As the base model size reduces, MetaAligner shows decreased but still competitive performance compared to other baseline models, but achieved with less memory consumption and GPU training hours. Impressively, the MetaAligner-1.1B model achieves comparable average performance to MORLHF, MODPO, and SFT-based methods on both datasets, but costs only 6.37%-48.71% of their GPU training hours, with a 6.36$\times$ smaller size than the LLaMA2-Chat-7B policy model. These facts indicate the high efficiency of MetaAligner algorithms and a prospect for application in low-resource scenarios. Compared to previous methods, MetaAligner models can also achieve balanced and stable performances in objective-wise evaluations, including contradictory objectives such as "Harmless" and "Helpful", without requiring explicit hyper-parameter tuning for achieving Pareto optimal solutions [11, 36, 20]. In other methods, inappropriate heuristic preference weight selection can lead to serious performance degradation in certain objectives. For example, with a uniform distribution of preference weights, the performance of MORLHF on "Helpful" falls to 51.2%, a huge gap to other methods. Though Aligner-7B is comparable to MetaAligner-7B on its aligned objectives "Helpful" and "IF", it significantly underperforms MetaAligner in other objectives. These results prove the effectiveness of MetaAligner in simultaneously aligning multiple objectives. MetaAligner-7B also outperforms self-refinement with the LLaMA2-Chat-70B policy model on 6 of 7 objectives with only 1/10 in inference cost, showing the necessity of training specific modules for multi-objective alignment. Ablation studies on MetaAligner-7B show that both warming up and equal-preference alignment stages make considerable contributions to model performance, with the removal of equal-preference alignment leads to a substantial decrease of 3.53% on HH-RLHF and 3.51% on UltraFeedback in average win rates.
Figure 3: Zero-shot alignment on 6 unseen objectives. In the x-axis, "Aligned Obj." denotes the 4 supervised objectives ($\diamond$" markers), and "+" denotes further addition of an unseen objective ($\circ$" markers). $\star$" denotes the win rates for the unseen objectives before all zero-shot alignments, "-." lines identify win rate fluctuations before alignment, and solid lines identify fluctuations after alignment.
4.4 Generalizable Alignment to Unseen Objectives
In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with six unseen objectives: "Specific", "Factual", "Readable", "Fair", "Repeat", and "Length" [10]. More details about these objectives are in Appendix C. We randomly select 2,700 queries from the UltraFeedback dataset and re-align the LLaMA2-Chat-70B outputs with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one, with 10 aligned objectives in total. Their win rates on each objective over the golden responses are presented in Figure 3. We have the following conclusions:
Performance on aligned objectives is maintained with additional unseen alignment objectives. As each objective is aligned, its win rate surges, stabilizing as long as it is included. On simultaneously aligning 10 objectives, MetaAligner-7B outperforms LLaMA2-Chat-70B outputs by an average of 14.25% on unseen objectives. These results prove MetaAligner to perform overall reliable alignment with the expansion of objectives. However, enhancements in one objective can affect performance in certain objectives due to their controversial nature, which is known as the "alignment tax" [11]. For example, aligning on "Fair" (+Fair) with MetaAligner-(7B, 13B) benefits its win rates, but harms performance on objectives such as "Readable" and "Factual" compared to when "Fair" is unaligned.
MetaAligner’s generalizability shows scalability. Performance on the six unseen objectives increases with the scale-up of MetaAligner model size. MetaAligner-1.1B provides limited improvement on most unseen objectives, but MetaAligner-7B extends the win rates to an average of 48.5%, and MetaAligner-13B further reaches 61.25%. MetaAligner-13B also more effectively aligns objectives such as "Length", where smaller models perform badly. This scalability is attributed to larger foundation models’ growing in-context learning ability, which enables accurate interpretations of the objective descriptions and instructions. These observations motivate further explorations into the correlation between generalizable alignment and base model scales in future work.
@@ 4.5 Evaluations of Objective-Wise Alignment
We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives. We utilize MetaAligner to perform six levels of alignments: unaligned, aligning on each objective ("Instruction following", "Helpful", "Honest", and "Truthful"), and full alignment. We leverage GPT-4 to score the responses ranging from 0 to 10. The results are shown in Figure 4. Experimental details and more results are shown in Appendix F. We have the following observations:
Objective-wise alignment improves performance on the primary target and boosts the performance on other objectives. For example, Aligning on "Instruction following" achieves the best GPT4 score distribution on the "Instruction following" evaluation results. It also significantly increases GPT4 scores on "Helpful", "Honest", and "Truthful" over the unaligned responses. This tendency holds with other policy models and alignment objectives. These results further prove the complex interplay among objectives, where correlations and contradictions [11] co-exist.